### v1

# Space-Time Bridge-Diffusion

Abstract
In this study, we introduce a novel method for generating new synthetic samples that are independent and identically distributed (i.i.d.) from high-dimensional real-valued probability distributions, as defined implicitly by a set of Ground Truth (GT) samples. Central to our method is the integration of space-time mixing strategies that extend across temporal and spatial dimensions. Our methodology is underpinned by three interrelated stochastic processes designed to enable optimal transport from an easily tractable initial probability distribution to the target distribution represented by the GT samples: (a) linear processes incorporating space-time mixing that yield Gaussian conditional probability densities, (b) their bridge-diffusion analogs that are conditioned to the initial and final state vectors, and (c) nonlinear stochastic processes refined through score-matching techniques. The crux of our training regime involves fine-tuning the nonlinear model, and potentially the linear models – to align closely with the GT data. We validate the efficacy of our space-time diffusion approach with numerical experiments, laying the groundwork for more extensive future theory and experiments to fully authenticate the method, particularly providing a more efficient (possibly simulation-free) inference.
keywords:
Optimal Transport, Diffusion Bridges, Stochastic Differential Equations, Fokker-Planck Equations, Generative Models, Sampling
thanks: We acknowledge use of ACCESS/NSF cloud computing resources via Discover grant Innovation in Generative Diffusion Modeling, MTH240001.
1 Introduction
This work addresses the quintessential inference challenge of generative learning: how to create new synthetic samples that are indistinguishable from a set of Ground Truth (GT) samples, assumed to be independently and identically distributed (i.i.d.) from an unknown probability distribution. Recently, diffusion models, leveraging Stochastic Differential Equations (SODE) to transform a tractable probability distribution into one that mirrors the GT distribution—thereby embodying the principles of optimal transport—have revolutionized this domain, surpassing competing methodologies Sohl-Dickstein et al. (2015); Ho et al. (2020); Song et al. (2021). Despite their empirical success, a comprehensive theoretical framework for diffusion modeling remains elusive. Our contribution lies in the development of such a framework, enhancing understanding while also introducing innovative learning and inference algorithms.
The essence of the diffusion model lies in its strategy to transform samples over an artificial timeline, adjusting phase-space mixing to achieve both repulsion among distinct samples and convergence towards Ground Truth (GT) samples, often in separate phases of the dynamics. This process, as practiced currently, unfolds in two stages. Initially, we introduce a straightforward stochastic process that indiscriminately mixes all elements of the high-dimensional state vector in a uniform manner, independent of the samples. This process is deemed tractable as its marginal probability density can be computed at any time without the need for simulation. Subsequently, the second stage involves the construction and training of a more intricate stochastic process that not only is influenced by the GT samples but also incorporates both temporal and spatial (inter-component) mixing which is nonlinear in the state variable. The nonlinearity is represented via score-matching- (or just score-) function represented via a Deep Neural Network (DNN).
The core contributions of this manuscript are grounded in the innovative proposal to decouple the traditional sequential relationship inherent in diffusion model construction. We present a methodology that embeds spatio-temporal mixing from the outset – directly within the base process – with possible subsequent refinement through optimization based on Ground Truth (GT) samples. A distinctive aspect of our work is parallel exploration of two options for the second stage process: one constructed forward in time, in harmony with the primary process’s temporal direction, and another fashioned in reverse time. These theoretical suggestions are bolstered by preliminary experiments, with additional validation efforts to be completed by the time of manuscript acceptance.
We structure the manuscript as follows. In Section 2, we introduce the foundational concepts of affine (linear in space) spatio-temporal bridge diffusion, setting the stage for more complex models and discussions. Section 3 advances into the nonlinear aspects of space-time bridge-diffusion, facilitated by score-matching functions parameterized through Deep Neural Networks (DNNs), beginning with a concise review of relevant literature on score-matching in both diffusion and bridge-diffusion processes. We then delineate two distinct bridge-diffusion approaches: a non-denoising scheme operating forward in time in Section 3.1, and a denoising scheme in reverse time in Section 3.2. Section 4 delves into optimizing these bridge-diffusion models to enhance alignment with Ground Truth (GT) data, subdivided into an exploration of affine space-time drift optimization in Section 4.1 and a heuristic approach for deriving GT-dependent drift from DNN-parameterized score-functions in Section 4.2. Our preliminary experimental findings, validating the theoretical frameworks discussed, are presented in Section 5. Concluding remarks and future directions are outlined in Section 6. The appendices provide in-depth technical details, with basic derivations of bridge diffusion processes based on Doob’s h-transform in Section A, and discussions on the general normal stochastic processes with spatial-temporal drift and their bridge diffusion counterparts in Sections B, C, and D, offering insights into their dual derivations.
2 Basic and Bridge Diffusion Processes with Affine Space-Time Mixing Drift
We examine a class of stochastic Bridge Diffusion (BD) processes for $t\in[0,1]$ where each state vector ${\bm{x}}(t)$ resides in $\mathbb{R}^{k}$, with $k$ representing, for example, equal the number of pixels in gray image. The foundational structure of a BD process, elaborated in Appendix A, originates from a primary stochastic model governed by Eq. (16). This model incorporates a deterministic drift vector ${\bm{f}}(t;{\bm{x}}(t))\in\mathbb{R}^{k}$ and a Wiener-diffusion matrix ${\bm{\kappa}}(t;{\bm{x}}(t))\in\mathbb{R}^{k\times k}$, which is positive-definite and $k\times k$-dimensional. The process undergoes further refinement by introducing an adjustment to the drift component, ensuring the emergence of a BD process that effectively acts as a conduit between two fixed points—transitioning from an initial state ${\bm{x}}(0)$ to a terminal state ${\bm{x}}(1)$, each represented by $\delta$-functions.
In this work, we focus on a particular instance of the basic stochastic process characterized by an affine drift, $\bm{f}(t;\bm{x}(t))\to\underline{\bm{A}}(t)\bm{x}(t)$, where $\underline{\bm{A}}(t)\in\mathbb{R}^{k\times k}$ denotes a time-dependent, positive-definite matrix independent of $\bm{x}(t)$. Correspondingly, we simplify the Wiener diffusion matrix to $\bm{\kappa}(t;\bm{x}(t))\to\bm{\kappa}(t)\in\mathbb{R}^{d\times d}$, leading to the following Stochastic Ordinary Differential Equation (SODE):
$\forall t\in[0,1]:\quad d\bm{x}(t)$ $=\underline{\bm{A}}(t)\bm{x}(t)dt+d\bm{W}(t),$ (1)
$\forall i,j:\ \mathbb{E}[dW_{i}(t)dW_{j}(t)]$ $=\kappa_{ij}(\tau)dt,$ (2)
with $\bm{W}(t)$ representing the Itô-regularized Wiener process, initiating the dynamics at $\bm{x}(0)$. This model is parameterized by $\bm{\psi}\vcentcolon=\{\underline{\bm{A}}(0\to 1),\bm{\kappa}(0\to 1)\}$. Furthermore, the model may commence from any $\bm{x}(t^{\prime})$ for $t^{\prime}\in[0,1]$, proceeding for $t\geq t^{\prime}$ until $t^{\prime\prime}>t^{\prime}$. While this scenario involves parameters $\{\underline{\bm{A}}(t^{\prime}\to t^{\prime\prime}),\bm{\kappa}(t^{\prime}\to t^{\prime\prime})\}$, we consistently employ the notation $\bm{\psi}$ for simplicity, sidestepping additional complexity.
The linearity of SODE as described by Eq. (1) ensures that the solution’s statistics, ${\bm{x}}(t)$, given an initial value ${\bm{x}}(0)$, adhere to a Gaussian distribution:
$\forall t,t^{\prime}\in[0,1];\ t\geq t^{\prime}:\quad p_{\bm{\psi}}\left({\bm{x}}(t)|{\bm{x}}(t^{\prime})\right)$ (3)
[…]
where $\underline{\bm{\Omega}}(t;\tau)$ represents a deterministic matrix process. This process is intricately linked to $\underline{\bm{A}}(\cdot)$ as defined by the following Ordinary Differential Equations (ODEs):
$0\leq\tau\leq t\leq 1:\quad\frac{d}{dt}\underline{\bm{\Omega}}(t;\tau)=\underline{\bm{A}}(t)\underline{\bm{\Omega}}(t;\tau),$ (4)
$\frac{d}{d\tau}\underline{\bm{\Omega}}(t;\tau)=-\underline{\bm{\Omega}}(t;\tau)\underline{\bm{A}}(\tau),\quad\underline{\bm{\Omega}}(\tau;\tau)={\bm{I}},$
with ${\bm{I}}$ denoting the $d\times d$ identity matrix.
We establish the following theorem:
Theorem 1
The BD adaptation of the fundamental stochastic process, as dictated by the SODE (22), is characterized by:
$t\in[0,1]:\quad d{\bm{x}}(t)=\bar{\bm{A}}(t)\left({\bm{x}}(t)-{\bm{x}}(1)\right)dt+d{\bm{W}}(t),$ (5)
where $\bar{\bm{A}}(\cdot)$ is formulated in terms of $\underline{\bm{A}}(\cdot)$ via implicit relationships:
$\bar{\bm{A}}(t)$ :$\vcentcolon=\frac{d}{dt}\log\left(\underline{\bm{\Omega}}^{-1}(1;t)-{\bm{I}}\right),$ (6)
$\frac{d}{dt}\underline{\bm{\Omega}}^{-1}(1;t)$ $=\underline{\bm{A}}(t)\underline{\bm{\Omega}}^{-1}(1;t),\quad\underline{\bm{\Omega}}^{-1}(1;1)={\bm{I}}.$
Remark 2
The deterministic linear ODE for $\underline{\bm{\Omega}}^{-1}(1;t)$, as outlined, aligns with the framework established in Eqs. (4). This can also be perceived through discrete-time relations as $t\to\Delta n/N,n=0,\ldots,N$, where the evolution of $\underline{\bm{\Omega}}(1;\Delta n/N)$ and its inverse are defined in the limit $\Delta\to 0$.
{pf}
The essence of the proof lies in employing Doob’s h-transform to evolve a basic SODE (1) into a BD SODE and juxtaposing it against the BD formulation (5). Detailed across Appendices A, B, C, and D, the proof begins with a generalization of the stochastic process governed by Eq. (22), extending it to both Eqs. (1) and (5). This generalization undergoes the BD transformation procedure, leading to an equation of special form akin to Eq. (5). Subsequent steps involve the adaptation of parameters to $\underline{\bm{A}}$ and ${\bm{0}}$, culminating in the derivation of Eq. (6) that implicitly connects $\underline{\bm{A}}$ and $\bar{\bm{A}}$. The conclusive part of the proof, which reconciles these relations, is elaborated in Appendix D.
Corollary 3
The statistics of the BD process (5) is given by the Gaussian distribution:
$p_{\bm{\psi}}\left({\bm{x}}(t)|{\bm{x}}(0);{\bm{x}}(1)\right)=\mathcal{N}\Bigg{(}{\bm{x}}(t)\Bigg{|}\bar{\bm{\Omega}}(t;0){\bm{x}}(0)+$ (7)
$\left({\bm{I}}-\bar{\bm{\Omega}}(t;0)\right){\bm{x}}(1);\int\limits_{0}^{t}dt^{\prime}\bar{\bm{\Omega}}(t;t^{\prime}){\bm{\kappa}}(t^{\prime})\left(\bar{\bm{\Omega}}(t;t^{\prime})\right)^{T}\Bigg{)},$
where the mean and covariance implicitly depend on the parameter set ${\bm{\psi}}\vcentcolon=\{\underline{\bm{A}}(0\to 1);{\bm{\kappa}}(0\to 1)\}$, as delineated by Eqs. (6).
3 Bridge-Diffusion with Nonlinear, Space-Time Mixing Drift
In this section, we build upon the foundation laid in the preceding discussion, which centered around linear SODEs with affine drift, particularly leveraging the insights from Theorem 1, to extend our modeling framework into the realm of nonlinear SODEs. This extension is pursued through two distinct methodologies, both underpinned by a Deep Neural Network (DNN) enforced score-matching strategy. This approach, initially proposed by Hyvärinen (2005) and further elaborated in Vincent (2011), has been adapted into discrete-time models in Sohl-Dickstein et al. (2015); Ho et al. (2020), and subsequently evolved into continuous-time diffusion processes in association with SODEs in Song et al. (2021). Remarkably, the continuous-time SODE-based algorithms introduced in Song et al. (2021) align with the earlier theoretical frameworks for denoising reverse time processes, such as those presented in Anderson (1982). However, our discussion begins by examining the non-denoising forward time approach, thus generalizing the work of Peluchetti (2021).
3.1 (Non-Denoising) Forward Time
The methodology presented in Peluchetti (2021) focuses on the special case of the Brownian bridge. We generalize this framework to the basic SODE (1) allowing a Gaussian probability density that conditions on the initial value ${\bm{x}}(0)$.
In a data-driven scenario, we assume the initial data ${\bm{x}}(0)$ is independently and identically distributed (i.i.d.) from a sample-tractable distribution $p_{0}(\cdot)$, such as one with independent components, $p_{0}({\bm{x}}(0))=\prod_{i=1}^{k}p_{0;i}(x_{i})$. Conversely, the final data ${\bm{x}}(1)$ is derived i.i.d. from the Ground Truth (GT) dataset ${\bm{X}}^{(S)}=({\bm{x}}^{(s)}|s=1,\cdots,S)$.
This approach, underpinned by Corollary 9 from Appendix A, employs a revised version of Eq. (21):
$t\in[0\to 1]:\quad d{\bm{x}}(t)=\Big{(}\underline{\bm{A}}(t){\bm{x}}(t)+$ (8)
${\bm{\kappa}}(t){\bm{s}}^{(f)}_{\bm{\theta}}(t;{\bm{x}}(t);{\bm{x}}(1))\Big{)}dt+d{\bm{W}}(t).$
where ${\bm{s}}^{(f)}_{\bm{\theta}}(t;{\bm{x}}(t);{\bm{x}}(1))$ denotes the vector score function, parameterized by a DNN and parameter set ${\bm{\theta}}$. The learning of this function is achieved by solving the optimization problem:
$\min_{\bm{\theta}}\mathbb{E}_{\text{FT}}\Big{[}\!\lambda(t)\Big{|}{\bm{s}}^{(f)}_{\bm{\theta}}(t;{\bm{x}}(t);{\bm{x}}(1))-$ (9)
$\nabla_{{\bm{x}}(t)}\log\left(p_{\bm{\psi}}({\bm{x}}(1)|{\bm{x}}(t))\right)\Big{|}^{2}\!\Big{]},\ \text{FT}\vcentcolon=\Big{\{}t\sim U([0,1]);$
${\bm{x}}(0)\sim p_{0}(\cdot);{\bm{x}}(1)\sim{\bm{X}}^{(S)};{\bm{x}}(t)\sim p_{\bm{\psi}}(\cdot|{\bm{x}}(0);{\bm{x}}(1))\Big{\}},$
with the expectation taken over the forward trajectory FT, defined by a set of distributions and sampling strategies for ${\bm{x}}(0)$, ${\bm{x}}(1)$, and ${\bm{x}}(t)$. The weight function $\lambda(t)$ serves as a hyperparameter, typically tuned empirically to enhance model performance.
Upon training the score function to optimal parameters ${\bm{\theta}}^{*}$, as determined by Eq. (9), Eq. (8) is then employed for inference, enabling the generation of new samples at $t=1$.
3.2 (Denoising) Reverse Time
In this subsection, we expand upon the denoising reverse time bridge-diffusion methodology as outlined in recent works Zhou et al. (2023); Li et al. (2023), focusing on scenarios where the affine-drift matrix $\underline{\bm{A}}(t)$ simplifies to a scalar coefficient $\alpha(t)$ multiplied by the identity matrix ${\bm{I}}$.
We invert the data-driven setting described previously, selecting initial data from the Ground Truth (GT) dataset $({\bm{x}}(0)\sim{\bm{X}}^{(S)})$, while the final data ${\bm{x}}(1)$ is drawn from a tractable distribution $p_{0}(\cdot)$. Under this configuration, the reverse time adaptations of Eqs. (8) and (11) are specified as follows, omitting detailed justifications which are thoroughly covered within the generative diffusion literature (e.g., see the review in Yang et al. (2022)):
$t\in[1\to 0]:\ d{\bm{x}}(t)=\Big{(}\bar{\bm{A}}(t)({\bm{x}}(t)-{\bm{x}}(0))$ (10)
$\hskip 79.6678pt-{\bm{\kappa}}(t){\bm{s}}^{(r)}_{\bm{\phi}}(t;{\bm{x}}(t);{\bm{x}}(0))\Big{)}dt+d\tilde{\bm{W}}(t),$
$\min_{\bm{\phi}}\mathbb{E}_{\text{RT}}\Big{[}\lambda(t)\Big{|}{\bm{s}}^{(r)}_{\bm{\phi}}(t;{\bm{x}}(t);{\bm{x}}(0))$ (11)
$-\nabla_{{\bm{x}}(t)}\log\left(p_{\bm{\psi}}({\bm{x}}(t)|{\bm{x}}(0);{\bm{x}}(1))\right)\Big{|}^{2}\Big{]},$
$\text{RT}=\Big{\{}t\sim U([0,1]);{\bm{x}}(0)\sim{\bm{X}}^{(S)};$
${\bm{x}}(1)\sim p_{0}(\cdot);{\bm{x}}(t)\sim p_{\bm{\psi}}(\cdot|{\bm{x}}(0);{\bm{x}}(1))\Big{\}},$
where $\tilde{\bm{W}}(t)$ denotes a standard Wiener process in reverse time with statistical properties analogous to ${\bm{W}}(t)$ from the forward time Eqs. (1) and (5). The probability $p_{\bm{\psi}}({\bm{x}}(t)|{\bm{x}}(0);{\bm{x}}(1))$ leverages the explicit formulation provided by Eq. (7), and ${\bm{s}}^{(r)}_{\bm{\phi}}(t;{\bm{x}}(t);{\bm{x}}(0))$ represents the reverse vector score function, parameterized by a DNN with parameters ${\bm{\phi}}$.
4 Optimal Affine Bridge-Diffusion
The generative models discussed in Section 2, which incorporate nonlinear spatio-temporal drifts learned from GT data via DNN, demonstrate high expressiveness. However, their effectiveness can vary with the GT data and is inherently tied to the underlying basic Gaussian process, as delineated by SODE (1) or its Bridge Diffusion (BD) counterpart (5). In Section 4.1, we examine the flexibility in selecting $\underline{\bm{A}}(\cdot)$ and/or $\bar{\bm{A}}(\cdot)$, utilizing the concept of ELBO (Evidence Lower BOund) optimality. Subsequently, Section 4.2 presents a heuristic approach for deriving a GT-dependent affine drift. This method leverages an iterative process, employing a Taylor expansion of a series of spatially nonlinear score-functions, to enhance model adaptability to GT data.
4.1 ELBO-Optimal Affine Bridge Diffusion
Definition 4 (max-FT-ELBO)
The Evidence Lower BOund (ELBO) Objective associated with Bridge Diffusion (BD) processes, which map initial to final samples, is defined as:
${\mathcal{O}}$ ($\left(\{\underline{A}(0\to 1)\};\{{\bm{\kappa}}(0\to 1)\};p_{0}(\cdot)\right)=$ (12)
[…]
with an implicit dependence on ${\bm{\psi}}=\{\underline{\bm{A}}(0\to 1),{\bm{\kappa}}(0\to 1)\}$, as influenced by Eqs. (7), (6), and (4), as well as the initial distribution $p_{0}(\cdot)$. The max-ELBO optimization seeks the optimal affine-drift and diffusion matrix values that maximize this sample-dependent objective (12):
($\left(\{\underline{A}^{*}(0\to 1)\};\{{\bm{\kappa}}^{*}(0\to 1)\};p_{0}^{*}(\cdot)\right)$ (13)
[…]
Remark 5
Several insights are provided to clarify the significance and nuances of the ELBO objective and its optimal parametrization, along with potential algorithmic implications:
  1. 1.
    Adhering to variational inference terminology Kingma and Welling (2022), we refer to the objective in Eq. (12) as ELBO to highlight it as the empirical expectation of the lower bound on the finite-time distribution of ${\bm{x}}(1)$, implicitly represented through the GT sample set ${\bm{X}}^{(S)}$.
  2. 2.
    The framework allows for an extension where $\underline{\bm{A}}(0\to 1)$ may also depend on ${\bm{x}}(1)$, resulting in $\underline{\bm{A}}(0\to 1;{\bm{x}}(1))$, a matrix function parameterizable via a DNN. This generalization can extend to the diffusion matrix as $\{{\bm{\kappa}}(0\to 1;{\bm{x}}(1))\}$.
  3. 3.
    For evaluation efficiency, it might be beneficial to streamline the parameter space in Eq. (13). Options include working with time-independent or co-dimensional matrices $\underline{\bm{A}}(0\to 1)$, restricting the class of initial distributions to those that are tractable, or tailoring the spatial structures of the affine-drift and diffusion matrices to specific applications, such as employing graph-Laplacian structures for image-related tasks.
The following statement follows from Theorem 1 by construction:
Corollary 6
For a class of probability distributions which can be represented by the mapping from $t=0$ of a tractable initial distribution, $p_{0}({\bm{x}}_{0})$ – via the basic stochastic affine-drift- diffusion process, governed by Eq. (5) with a particular $\underline{\bm{A}}(0\to 1)$ and ${\bm{\kappa}}(0\to 1)$ – to $p({\bm{x}}(1))$, the following statements hold:
  1. 1.
    The BD version of the basic mapping procedure – where the basic process Eq. (5) is replaced by its BD version (1), with $\bar{\bm{A}}(0\to 1)$ expressed via $\underline{\bm{A}}(0\to 1)$ according to Eqs. (4,6) and with samples from $p({\bm{x}}(1))$ used as the pined values at $t=1$ for $S$ samples – will converge to the basic mapping procedure in the $S\to\infty$ limit.
  2. 2.
    Unless a degeneracy, the max-ELBO optimization (13) will result in the optimal values ($\Big{(}\{\underline{A}^{*}(0\to 1)\}$; $\{{\bm{\kappa}}^{*}(0\to 1)\};p_{0}^{*}(\cdot)\Big{)}$ which converges to the true values in the $S\to\infty$ limit.
We now propose a bold working hypothesis, pending verification through numerical experiments:
Conjecture 7
Initiating with samples ${\bm{x}}(0)$ drawn from $p_{0}(\cdot)$ and evolving these samples in accordance with the SODE (1), utilizing optimal affine-drift and diffusion matrices, alongside an initial distribution that satisfies the max-ELBO optimization criterion (13), we hypothesize that the resulting samples ${\bm{x}}^{(s)}(1)$ will, in the $S\to\infty$ limit, become i.i.d. according to a probability distribution that either represents a general position or belongs to a sufficiently large and practically relevant subclass thereof, as delineated by the ground truth sample set ${\bm{X}}^{(S)}$.
4.2 Affine Optimality via (Nonlinear) Score-Function
In this subsection, we explore how the use of a DNN-parameterized score function for the reverse process, as delineated in Eq. (11), provides heuristic alternatives to the direct resolution of Eq. (13). While an analogous procedure could be applied to the forward process described by Eq. (9), we omit it here for brevity.
Suppose the optimal score function ${\bm{s}}^{(r)}_{{\bm{\phi}}^{*}}({\bm{x}}(t);{\bm{x}}(0))$ has been identified. In that case, the optimal affine drift can be derived by approximating this score function with the leading term of its Taylor expansion around the deviation of ${\bm{x}}(t)$ from ${\bm{x}}(1)$:
$s^{(r)}_{i;{\bm{\phi}}^{*}}(t;{\bm{x}}(t);{\bm{x}}(0))\approx\sum_{j}(x_{j}(t)-x_{j}(0))H^{(r)}_{ij}(t;{\bm{x}}(0)),$ (14)
$H^{(r)}_{ij}(t;{\bm{x}}(0))=\partial_{x_{j}(t)}s^{(r)}_{j;{\bm{\phi}}^{*}}({\bm{x}}(t);{\bm{x}}(0))\Big{|}_{{\bm{x}}(t)={\bm{x}}(0)}.$
We can deduce the right-hand side of Eq. (14) from the optimal score function, represented by a DNN, using differential programming techniques, such as those discussed in Meng et al. (2021). Alternatively, this can be accomplished by addressing a "spatially-smooth" version of Eq. (11) as follows:
$\min_{\bm{\theta}}$ $\mathbb{E}_{\text{RT}}\Bigg{[}\lambda(t)\sum_{i,j}\Big{(}H^{(r)}_{ij;\bm{\theta}}(t;{\bm{x}}(0))-$ (15)
$\nabla_{x_{i}(t)}\nabla_{x_{j}(t)}\log\left(p_{\bm{\psi}}({\bm{x}}(t)|{\bm{x}}(0);{\bm{x}}(1))\right)\Big{|}_{{\bm{x}}(t)={\bm{x}}(0)}\Big{)}^{2}\Bigg{]},$
thereby estimating the sample-averaged Hessian of the log-probability.
Solving optimization (15) yields the Hessian, facilitating a simulation-free approach to inference as outlined in the following pseudo-algorithm for generating new samples, ${\bm{x}}^{(\text{new})}$:
  1. 1.
    Update $\underline{\bm{A}}(t;{\bm{x}}(0))\leftarrow\underline{\bm{A}}(t)+H^{(r)}_{{\bm{\theta}}^{*}}(t;{\bm{x}}(0))$, consequently updating ${\bm{\psi}}$ to be dependent on ${\bm{x}}(0)$.
  2. 2.
    Sample ${\bm{x}}(0)\sim{\bm{X}}^{(S)}$.
  3. 3.
    Generate ${\bm{x}}^{(\text{new})}\sim p_{\bm{\theta}}({\bm{x}}^{(\text{new})}|{\bm{x}}(0);{\bm{x}}(1))$.
This process – iterative learning of the score function with respect to the current ${\bm{\psi}}$ (now dependent on ${\bm{x}}(0)$) and subsequent updating of ${\bm{\psi}}$ – can be repeated until convergence, or at least as long as the ELBO objective in (13) shows improvement.
5 Numerical Experiments
In our investigation, we undertake numerical experiments utilizing the MNIST dataset, an established benchmark in the machine learning domain for handwritten digit recognition, which consists of 70,000 images with dimensions of $28\times 28$ pixels. For the purpose of training our model, we allocate 60,000 images, reserving the remaining 10,000 images to evaluate the model’s capability through the Fréchet Inception Distance (FID) metric. This metric assesses the quality and diversity of generated images by comparing their statistical properties in the feature space of a pre-trained Deep Neural Network (DNN) against those of real images. Importantly, a lower FID score signifies a closer approximation to the actual data distribution, reflecting the model’s enhanced ability to produce images that are perceived as more realistic by human evaluators. This metric is pivotal for our analysis as it provides a robust indicator of image quality that aligns well with human visual judgment.
For our Deep Neural Network (DNN), we opt for the NCSN++ architecture [^1] https://github.com/yang-song/score_sde_pytorch, recognized for its effectiveness in generative modeling. The training is conducted on an NVIDIA A100 GPU, spanning over 100 epochs to ensure sufficient learning and adaptation of the model. We utilize a batch size of 64, employing the Adam optimizer for its efficiency in handling sparse gradients and adaptability in large-scale problems, with a learning rate set at $10^{-4}$. This configuration is chosen to balance the computational efficiency and the convergence rate, optimizing the model’s performance in generating high-quality images.
Our investigations extend to stochastic Gaussian processes governed by basic Stochastic Ordinary Differential Equations (SODE), outlined in Eq. (1). Additionally, we explore their Bridge-Diffusion (BD) variants, delineated by Eq. (5), alongside non-Gaussian stochastic processes operating in both forward and reverse time. These processes are integral to generative models trained on Ground Truth (GT) data using Deep Neural Networks (DNNs), as elaborated in Sections 3.1 and 3.2.
We choose the "temporal" model, often referred to as the Brownian Bridge, for the benchmark. This involves BD matrix and covariance matrix which are co-dimensional to the unit matrix: $\bar{\bm{A}}(t)=\bm{I}/(1-t)$ and $\bm{\kappa}(t)=\bm{I}$. As derived from Eqs. (7), this model yields a Gaussian process characterized by the mean vector, ${\bm{\mu}}(t)=(1-t)\bm{x}(0)+t\bm{x}(1)$, and the covariance matrix, $\bm{\Sigma}(t)=t(1-t)\bm{I}$. The FID scores for generative model, tailored to this temporal BD model and applied to MNIST GT data, are depicted in Fig. 1. This figure facilitates a comparison between the forward time scheme (discussed in Section 3.1) and the reverse time scheme (outlined in Section 3.2), revealing that for fewer than 1000 discretization steps, the forward time scheme is more effective. Conversely, the reverse time scheme surpasses the forward scheme in performance when the discretization steps exceed 1000.
Figure 1: Comparison of FID scores between forward time and reverse time generative models employing the Brownian Bridge scheme, characterized by $\bar{\bm{A}}(t)=\bm{I}/(1-t)$ and $\bm{\kappa}(t)=\bm{I}$. Notably, with 1000 discretization steps, both models achieve an FID score of approximately 2, indicating a high degree of similarity in image quality and diversity to the real data distribution at this level of discretization.
We adopt $\bar{\bm{A}}(t)={\bm{L}}/(1-t)$ and $\bm{\kappa}(t)=\bm{I}$ in our space-time mixing model, where ${\bm{L}}$ signifies the graph-Laplacian matrix that models interactions between nearest-neighbor pixels on a square grid. Defined as $\bm{L}=\bm{I}-\bm{D}^{-1/2}\bm{J}\bm{D}^{-1/2}$, where $\bm{J}$ is the adjacency matrix and $\bm{D}$ a diagonal matrix representing each pixel’s connectivity degree, $\bm{L}$ emerges as a symmetric positive semi-definite matrix. Its eigenvalues $\lambda_{1},\cdots,\lambda_{d}$, span from 00 to $2$, leading to $\bm{L}=\bm{P}\text{diag}(\lambda_{1},\cdots,\lambda_{n})\bm{P}^{\top}$ in the eigenvalue decomposition. To address the zero eigenvalue, a standard pseudo-inverse is used to resolve the formal singularity. The resultant space-time BD model, as per Eqs. (7), produces a Gaussian process with mean vector, $\bm{\mu}(t)=(1-t)^{\bm{L}}\bm{x}(0)+(\bm{I}-(1-t)^{\bm{L}})\bm{x}(1)$, and covariance matrix, $\bm{\Sigma}(t)=\bm{P}\ \text{diag}\left(\frac{(1-t)^{2\lambda_{1}}-(1-t)}{2\lambda_{1}-1},\cdots,\frac{(1-t)^{2\lambda_{n}}-(1-t)}{2\lambda_{n}-1}\right)\ \bm{P}^{T}$.
In our framework, each image is represented as a grid graph with dimensions $28\times 28$, leading to a Laplacian matrix of size $784\times 784$. This approach facilitates the exploration of the space-time BD generative modeling effectiveness, with FID scores presented in Fig. (2). Analysis reveals that for discretization steps under 1000, the forward time scheme outperforms the reverse time scheme in terms of FID score. Conversely, at 1000 discretization steps, the reverse time scheme demonstrates a marginal advantage, underscoring the nuanced performance dynamics between these schemes. Samples of generated images, with 1000 steps discretizing the $[0,1]$ time interval and using spatio-temporal bridge and reverse time, are shown in Fig. 3.
Figure 2: FID scores comparison for space-time BD models employing forward and reverse time approaches. This illustration underscores the models’ performance in generative tasks, as elaborated in the accompanying text. Notably, with 1000 discretization steps, the FID score approximates 1.5, highlighting the nuanced efficacy of these models in capturing the data’s underlying distribution with a high degree of fidelity.
Figure 3: This visualization showcases the model’s generative capability to synthesize high-fidelity images by integrating space time BD model within the reverse time framework and with 1000 discretization steps.
Comparing Fig. 1 with Fig. 2, we note that spatial-temporal mixing achieves superior FID scores using fewer discretization steps. However, this advantage requires evaluating the graph Laplacian indicating an added computational cost (for training and not inference). To extend this method to higher-resolution images, adopting techniques from auto-encoders, as exemplified in the Stable Diffusion (text-to-image generation) methodology Rombach et al. (2021), offers a viable pathway for scaling the approach effectively.
6 Conclusions and Path Forward
In this work, we introduce a novel spatio-temporal generalization of Bridge-Diffusion (BD) processes, a topic that has garnered considerable attention in recent AI research. Our contribution lies in expanding the class of tractable Gaussian processes within the BD framework by incorporating a novel spatial mixing that is linear/affine with respect to the state-vector, in addition to the temporal mixing found in existing literature. We further develop this affine spatio-temporal BD concept through the integration of more complex nonlinear dynamics, utilizing score-matching functions and Deep Neural Networks (DNNs). Our exploration leads to the proposal of innovative BD schemes, delineating the distinctions between non-denoising (forward in time) and denoising (reverse in time) approaches. Moreover, we investigate the optimization of these schemes to enhance alignment with Ground Truth (GT) data, presenting a theoretical framework for affine space-time drift optimization and the derivation of optimal and GT-dependent affine drifts. Through empirical validation, we partially substantiate our theoretical claims, concluding that the integration of spatial-temporal mixing significantly enhances the inference performance of score-based, nonlinear DNN-trained generative models. Our comparative analysis of forward time (without denoising) versus reverse time (with denoising) schemes reveals that the former yields superior results with a sufficiently small number of discretization steps, while both approaches exhibit comparable FID performances as the number of steps approaches 1000.
As we await feedback from the review process, our efforts will persist in the experimental validation of the theoretical constructs presented within our manuscript. We aim to delve deeper into the ELBO-based affine drift optimization scheme introduced in Section 4.1, as well as the iterative learning process for the score-function, as detailed in Section 4.2, applying these methodologies across both forward and reverse time DNN learning approaches for score functions. The refinement and implementation of the optimal affine scheme will further enable us to test Conjecture 7, opening a promising avenue toward simulation-free inference solutions. Additionally, we intend to investigate efficient representations for specific subclasses of affine BD drift terms, as discussed in Section 5. This exploration is anticipated to facilitate the extension of our numerical experiments to encompass larger-scale (higher resolution) images, thereby broadening the scope and applicability of our findings.
Appendix A General Bridge-Diffusion Processes
We commence this Appendix by revisiting the classic problem of the diffusion bridge, closely related to what is known as Doob’s h-transform. This topic has been thoroughly explored in the literature, including Chapter 7.5 of Särkkä and Solin’s book Särkkä and Solin (2019), as well as Chapter IV.6.39 in the comprehensive text by Rogers and Williams Rogers and Williams (2000). Moreover, it has recently been contextualized within the field of generative AI modeling, as discussed in Section 3.1 of Peluchetti et al. Peluchetti (2021).
We consider a general drift-diffusion process that governs the stochastic dynamics of a $d$-dimensional vector, denoted by ${\bm{x}}(\tau)\vcentcolon=(x_{i}(\tau)\in\mathbb{R}\,|\,i=1,\ldots,d)$:
$\tau\in[0,1]:\quad d{\bm{x}}(\tau)={\bm{f}}(\tau;{\bm{x}}(\tau))\,d\tau+d{\bm{W}}(\tau),$ (16)
$\forall\,i,j=1,\ldots,d:\ \mathbb{E}[dW_{i}(\tau)\,dW_{j}(\tau)]=\kappa_{ij}({\bm{x}}(\tau),\tau),$
where ${\bm{W}}(\tau)$ represents a Wiener process characterized by the space-time dependent covariance matrix ${\bm{\kappa}}({\bm{x}}(\tau),\tau)$. The conditional probability $p_{\bm{\psi}}{\bm{x}}(\tau)|{\bm{x}}(\tau^{\prime}))$, representing the probability of observing ${\bm{x}}$ at time $\tau$ given the observation of ${\bm{x}}^{\prime}$ at an earlier time $\tau^{\prime}$, where $01>\tau>\tau^{\prime}>$ 0, is governed by the Fokker-Planck (FP) equations for forward and reverse dynamics, respectively:
$\left(\partial_{\tau}-\hat{\mathcal{L}}^{*}_{\tau}({\bm{x}}(\tau))\right)p_{\bm{\psi}}{\bm{x}}(\tau)|{\bm{x}}(\tau^{\prime}))=0,$ (17)
^$\hat{\mathcal{L}}^{*}_{\tau}({\bm{x}})\equiv-\partial_{i}f_{i}(\tau;{\bm{x}})+\frac{1}{2}\partial_{i}\partial_{j}\kappa_{ij}({\bm{x}},\tau),$
$\left(\partial_{\tau}+\hat{\mathcal{L}}_{\tau}({\bm{x}}(\tau))\right)p_{\bm{\psi}}{\bm{x}}(1)|{\bm{x}}(\tau))=0,$ (18)
^$\hat{\mathcal{L}}_{\tau}({\bm{x}})\equiv f_{i}(\tau;{\bm{x}})\partial_{i}+\frac{1}{2}\kappa_{ij}({\bm{x}},\tau)\partial_{i}\partial_{j},$
where $\hat{\mathcal{L}}$ and $\hat{\mathcal{L}}^{*}$ denote the direct and adjoint generators of the stochastic dynamics, respectively. We adopt the notations and terminology from Särkkä and Solin Särkkä and Solin (2019), utilizing shorthand $\partial_{i}$ to denote $\partial_{x_{i}}$ and invoking the Einstein summation convention for repeated indices. The forward (17) and reverse (18) forms of the FP equation are well-established within stochastic calculus, most succinctly derived by marginalizing the discretized Feynman-Kac (path-integral) formulations for a joint probability distribution. This approach involves a sequence of Gaussian integrations, followed by taking the continuous-time limit.
Next, we define the transformed probability density
$\tilde{p}({\bm{x}}(\tau)|{\bm{x}}(\tau^{\prime}))\vcentcolon=p({\bm{x}}(\tau)|{\bm{x}}(\tau^{\prime}))\frac{p({\bm{x}}(1)|{\bm{x}}(\tau))}{p({\bm{x}}(1)|{\bm{x}}(\tau^{\prime}))},$ (19)
and note that: (a) it constitutes a legitimate probability density, as it is non-negative and integrates to one over the entire domain of ${\bm{x}}(\tau)$; and (b) as $\tau\to 1$, it converges to the Dirac delta function $\delta({\bm{x}}(\tau)-{\bm{x}}(1))$.
The remarkable statement is encapsulated in the following theorem:
Theorem 8 (Fokker-Planck for BD)
The transformed probability density $\tilde{p}({\bm{x}}(\tau)|{\bm{x}}(\tau^{\prime}))$, as defined in Eq. (19), satisfies the bridge diffusion Fokker-Planck equation:
$\Big{(}\partial_{\tau}-\hat{\mathcal{L}}^{*}_{\tau}({\bm{x}})$ (20)
$\hskip 14.22636pt+\partial_{{\bm{x}}_{i}(\tau)}\kappa_{ij}({\bm{x}}(\tau),\tau)s_{j;{\bm{x}}(1)}({\bm{x}}(\tau);\tau)\Big{)}\tilde{p}({\bm{x}}(\tau)|{\bm{x}}(\tau^{\prime}))=0,$
$\hskip 14.22636pts_{j;{\bm{x}}(1)}({\bm{x}}(\tau);\tau)\vcentcolon=\partial_{{\bm{x}}_{j}(\tau)}\log\left(p({\bm{x}}(1)|{\bm{x}}(\tau))\right).$
{pf}
The proof is algebraic and straightforward. We apply the operator from the left-hand side of Eq. (20) to the transformed probability density $\tilde{p}({\bm{x}}(\tau)|{\bm{x}}(\tau^{\prime}))$, as defined in Eq. (19). By employing the chain rule for differentiation and subsequently invoking Eqs. (17) and (18), we represent the temporal derivatives $\partial_{\tau}p({\bm{x}}(\tau)|{\bm{x}}(\tau^{\prime}))$ and $\partial_{\tau}p({\bm{x}}(1)|{\bm{x}}(\tau))$ in terms of the respective spatial derivatives of $p({\bm{x}}(\tau)|{\bm{x}}(\tau^{\prime}))$ and $p({\bm{x}}(1)|{\bm{x}}(\tau))$. Upon collecting all terms, they sum to zero, thus completing the proof.
The ensuing corollary is a direct consequence of Theorem 8 and is also aligned with Theorem 7.11 in Särkkä and Solin Särkkä and Solin (2019):
Corollary 9 (Stochastic Bridge Diffusion Process)
Given the basic SDE as in Eq. (16), we now aim to modify the process by conditioning the solution to satisfy ${\bm{x}}(1)$ at time $t=1$. Additionally, assuming that the conditional probability density $p({\bm{x}}(1)|{\bm{x}}(t))$, associated with the basic SDE (16) and defined in Eq. (17), has been computed as a function of ${\bm{x}}(\tau)$. It results in the following stochastic process:
$\tau\in[0,1]:\quad d{\bm{x}}(\tau)=\Big{(}{\bm{f}}(\tau;{\bm{x}}(\tau))$ (21)
$\hskip 14.22636pt+{\bm{\kappa}}({\bm{x}}(\tau),\tau)\nabla_{{\bm{x}}(\tau)}\log\left(p({\bm{x}}(1)|{\bm{x}}(\tau))\right)\Big{)}d\tau+d\bm{W}(\tau).$
Appendix B Normal Stochastic Process
Consider the special case of the stochastic drift-diffusion process (16) with affine drift, ${\bm{f}}({\bm{x}}(\tau);\tau)={\bm{A}}(\tau){\bm{x}}(\tau)+{\bm{c}}(\tau)$, and ${\bm{x}}(\tau)$ independent diffusion, ${\bm{\kappa}}({\bm{x}}(\tau);\tau)\to{\bm{\kappa}}(\tau)$. Then evaluation of the basic Stochastic Differential Equation (SODE) – with the initial condition at $t=0$ pined/fixed to ${\bm{x}}(0)={\bm{x}}_{0}$:
$\tau\in[0,1]:\ d{\bm{x}}(\tau)\!=\!\left({\bm{A}}(t){\bm{x}}(\tau)\!+\!{\bm{c}}(\tau)\right)d\tau+d{\bm{W}}(\tau).$ (22)
Statistics of ${\bm{x}}(\tau)$, conditioned to the initial value ${\bm{x}}(0)$ is Gaussian:
$p\left({\bm{x}}(t)|{\bm{x}}(\tau)\right)={\mathcal{N}}\Bigg{(}{\bm{x}}(t)\Bigg{|}{\bm{\Omega}}(t;\tau){\bm{x}}(\tau)$
$+\int\limits_{\tau}^{t}dt^{\prime}{\bm{\Omega}}(t;t^{\prime}){\bm{c}}(t^{\prime});\int\limits_{\tau}^{t}dt^{\prime}{\bm{\Omega}}(t;t^{\prime}){\bm{\kappa}}(t^{\prime}){\bm{\Omega}}(t;t^{\prime})\Bigg{)},$ (23)
$t^{\prime}\in[\tau,t]:\ \frac{d}{dt^{\prime}}{\bm{\Omega}}(t^{\prime};\tau)={\bm{A}}(t^{\prime}){\bm{\Omega}}(t^{\prime};\tau),$ (24)
$\frac{d}{d\tau}{\bm{\Omega}}(t^{\prime};\tau)=-{\bm{\Omega}}(t^{\prime};\tau){\bm{A}}(\tau),\quad{\bm{\Omega}}(\tau;\tau)={\bm{I}},$
where thus ${\bm{\Omega}}(t^{\prime};\tau)$ is a deterministic matrix process which depends implicitly on $({\bm{A}}(\tau\to t)$; and we shortcut notations omitting dependence of the probability density on $\underline{\bm{A}}(0\to 1)$ and ${\bm{\kappa}}(0\to 1)$.
Appendix C Normal Bridge-Diffusion Process
In this Appendix we combine ideas from Appendix B and Appendix A. The goal is to construct a Bridge-Diffusion process of a special form – Gaussian process, starting and resulting in $\delta$-function with a specific evolution of the mean vector and of the covariance matrix. Eq. (22) with ${\bm{A}}(t)$ is our starting point here. The conditional probability density of the normal process – with the initial condition at $\tau$ and the observation point $t$ – is
$p^{(\text{SEN})}_{\bm{\psi}}\!\!\left({\bm{x}}(t)|{\bm{x}}(\tau)\right)\!=\!{\mathcal{N}}\Big{(}{\bm{x}}(t)\Big{|}{\bm{\Omega}}(t;\tau){\bm{x}}(\tau)\!+\!\int\limits_{\tau}^{t}\!\!dt^{\prime}{\bm{\Omega}}(t;t^{\prime}){\bm{c}}(t^{\prime});$
${\bm{\Sigma}}(\tau)\Big{)},\quad{\bm{\Sigma}}(\tau)\vcentcolon=\int\limits_{\tau}^{t}dt^{\prime}{\bm{\Omega}}(t;t^{\prime}){\bm{\kappa}}(t^{\prime})\left({\bm{\Omega}}(t;t^{\prime})\right)^{T},$ (25)
where ${\bm{\Sigma}}(\tau)$ is the covariance matrix; and we shortcut notations omitting dependencies on ${\bm{A}}(\tau\to t;\tau)$; ${\bm{c}}(\tau\to t)$; and ${\bm{\kappa}}(\tau\to t)$.
Next, following the general construction of Appendix A we arrive at the Bridge Diffusion (BD) version of the process described by Eq. (25) with the BD pined at $\tau=t$ to ${\bm{x}}(t)$:
$d{\bm{x}}(\tau)=\Big{(}\!{\bm{A}}(\tau){\bm{x}}(\tau)+{\bm{c}}(\tau)$ (26)
$+{\bm{s}}\left({\bm{x}}(\tau);\tau;{\bm{x}}(t);\{{\bm{\theta}}(\tau\to t)\}\right)\Big{)}d\tau\!+d{\bm{W}}(\tau)$
${\bm{s}}\left({\bm{x}}(\tau);\tau;{\bm{x}}(t);\{{\bm{\theta}}(\tau\to t)\}\right)$ (27)
:$\vcentcolon=\nabla_{{\bm{x}}(\tau)}\log\left(p^{(\text{SEN})}\left({\bm{x}}(t)|{\bm{x}}(\tau);\{{\bm{\theta}}(\tau\to t)\}\right)\right)$
[…]
We observe that the BD-correction, that is the ${\bm{s}}$-term, results in “re-normalization" of the ${\bm{A}}(\cdot)$ and ${\bm{c}}(\cdot)$ (functions-) coefficients and can thus be viewed as a correction which allows to restate the BD-SEN Eq. (26) as a “renormalization" of the standard SEN process described by Eq. (22):
$d{\bm{x}}(\tau)$ $\!=\!\Big{(}\tilde{\bm{A}}(\tau){\bm{x}}(\tau)\!-\!\left({\bm{\Sigma}}(\tau)\right)^{-1}{\bm{x}}(t)\!+\!{\bm{\varsigma}}(\tau)\Big{)}d\tau+d{\bm{W}}(\tau),$
$\tilde{\bm{A}}(\tau)$ :$\vcentcolon={\bm{A}}(\tau)+\left({\bm{\Sigma}}(\tau)\right)^{-1}{\bm{\Omega}}(t;\tau),$
[…]
Then, the conditional probability density associated with Eq. (26) becomes:
$p^{(\text{SEN-BD})}\left({\bm{x}}(\tau)|{\bm{x}}(0);{\bm{x}}(t)\right)={\mathcal{N}}\Big{(}{\bm{x}}(\tau)\Big{|}\tilde{\bm{\Omega}}(\tau;0){\bm{x}}(0)$ (29)
$-\tilde{\bm{\Lambda}}(\tau){\bm{x}}(t)+\tilde{\bm{\varsigma}}(\tau);\tilde{\bm{\Sigma}}(\tau)\Big{.}\Big{)},\ \tilde{\bm{\Lambda}}(\tau)\!\vcentcolon=\!\!\int\limits_{0}^{\tau}\!\!dt^{\prime}\tilde{\bm{\Omega}}(\tau;t^{\prime})\left({\bm{\Sigma}}(t^{\prime})\right)^{-1},$
[…]
where $\tilde{\bm{\Omega}}(t^{\prime};\tau)$ is defined implicitly as a solution of the following matrix differential equation:
$\frac{d}{dt^{\prime}}\tilde{\bm{\Omega}}(t^{\prime};\tau)=\tilde{\bm{A}}(t^{\prime})\tilde{\bm{\Omega}}(t^{\prime};\tau),\quad\tilde{\bm{\Omega}}(\tau;\tau)={\bm{I}}.$ (30)
Derivations similar to ones presented in this Appendix were reported earlier in Barczy and Kern (2010).
Appendix D Normal Bridge Diffusion Process: Alternative Derivation
[…]
$\forall t\geq\tau,$ $t,\tau\in[0,1]:\ \frac{d}{dt}\bar{\bm{\Omega}}(t;\tau)=\bar{\bm{A}}(t)\bar{\bm{\Omega}}(t;\tau),$ (31)
$\frac{d}{d\tau}\bar{\bm{\Omega}}(t;\tau)=-\bar{\bm{\Omega}}(t;\tau)\bar{\bm{A}}(\tau),\quad\bar{\bm{\Omega}}(\tau;\tau)={\bm{I}},$
we arrive at the marginal probability distribution of ${\bm{x}}(t)$ conditioned to ${\bm{x}}(0)$ at $t=0$ and to ${\bm{x}}(1)$ at $t=1$ stated in Eq. (7).
We observe that if $\bar{\bm{A}}(t)$ is such that
$\forall t\in[0,1]:|\bar{\bm{A}}(t)|<+\infty,\ \&\ \bar{\bm{\Omega}}(1;t)\to 0\ \text{at}\ t\to 1,$ (32)
it makes the process described by Eq. (5) a Normal Bridge Diffusion (NBD) process.
Notice that constructing the NBD process described by Eqs. (7,32) we acted directly – bypassing the procedure described in Appendix C, where we first introduced a Normal Diffusion (ND) process and then applied to it the BD construction of Appendix A. It is thus natural to look for expressing $\bar{A}(t)$ in Eq. (5) via parameters of the ND process (1). Juxtaposing Eq. (5) term-by-term to Eq. (28) with ${\bm{A}}(\cdot),{\bm{\Sigma}}(\cdot)$ and ${\bm{\Omega}}(\cdot;\cdot)$ substituted respectively by $\underline{\bm{A}}(\cdot),\underline{\bm{\Sigma}}(\cdot)$ and $\underline{\bm{\Omega}}(\cdot;\cdot)$, we first of all observe that ${\bm{c}}(\cdot)$ can be set to ${\bm{0}}$, and then arrive at the relations:
$\underline{\bm{A}}(t)+\left(\underline{\bm{\Sigma}}(t)\right)^{-1}\underline{\bm{\Omega}}(1;t)=\bar{\bm{A}}(t),\ \left(\underline{\bm{\Sigma}}(t)\right)^{-1}=\bar{\bm{A}}(t).$ (33)
Resolving the relations, while also accounting for the "underline" version of Eq. (24) counted backwards in time we derive Eq. (6).